If you walked 2 3/4 miles on Monday and then walked 3 1/3 miles on Tuesday, how many miles did you walk during those two days?
Found 2 solutions by josgarithmetic, josmiceli:
Answer by josgarithmetic(39838)
Answer by josmiceli(19441)

2 and 3/4 is 2 and 9/12
-------------------------------
3 and 1/3 is 3 and 4/12
-------------------------------
Add the whole numbers and then the fractions
2 + 3 = 5
9/12 + 4/12 = 13/12
13/12 = 1 and 1/12
Now you have:
5 + 1 + 1/12 = 6 1/12